2. A 0.825-kg block of iron, with an average specific heat of 5.60 x102 J/kg K, is initially at a temperature of 352°C. It is placed in a calorimeter that has 40.0 g of water at 20.0°C. What is the final thermal equilibrium temperature? If your answer if 100.0°C, calculate the amount of water that remains in liquid form. Treat the mass and heat capacity of the calorimeter as neglible.
Answer #1

Let final temperature be T and all water has been evaporated, so we assume T>100,

heat lost by iron block = heat gained by water

miron*Siron* delta Tiron = mw*Sw* delta T water + mw L

0.825*560*[352-T] = 0.040*4186*[T-20] + 0.040*2260e3

T = 120 degree C answer